# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ 4.3 Asymptotic Normality of LOORA-DM
This section is focused on the analysis of the large-sample behavior of the leave-one-out regression-adjusted difference-in-means (LOORA-DM) estimator under complete random assignment. Our objective is to establish that in large samples, the random leave-one-out coefficients used in LOORA-DM converge to a deterministic population limit and that the estimator itself satisfies a central limit theorem with a well-defined asymptotic variance.
@@
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{d}}\big\|{\bm{X}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}_{-i}\big\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2}.$ (15)
Here $\widetilde{{\bm{y}}}^{(-i)}$ rescales the observed outcomes in a manner that depends on whether the $i$th unit is assigned to treatment or control. This rescaling is necessary because, under complete random assignment, the treatment indicators are not independent across units, and the weights must therefore be chosen to preserve unbiasedness of the estimator.
To study the limiting behavior of these regression adjustments, define the population analogue
$\boldsymbol{\beta}^{*}_{\lambda}:=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{d}}\left\|{\bm{X}}{\bm{b}}-\frac{\widetilde{{\boldsymbol{\mu}}}}{n}\right\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2},$ (16)
where $\widetilde{{\boldsymbol{\mu}}}=n_{C}{\bm{y}}^{(1)}+n_{T}{\bm{y}}^{(0)}$. This limit represents the deterministic coefficient vector that best captures the relation between covariates and the signal component $\widetilde{{\boldsymbol{\mu}}}/n$ in the population. The next lemma shows that the leave-one-out coefficients converge in probability to this population limit.
Lemma 6.
Suppose the potential outcomes, covariates, and treatment assignment satisfy Assumptions˜1, 2 and 3, and let $\lambda\geq 0$ be fixed. Then $\boldsymbol{\beta}^{*}_{\lambda}$, as defined in (16), exists. Moreover, for all $i\in\mathbb{N}$,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}.$
This result implies that the random fluctuations in the estimated adjustment coefficients vanish asymptotically. Hence, in large samples, the behavior of LOORA-DM is equivalent to that of an “oracle” estimator that uses the deterministic coefficient $\boldsymbol{\beta}^{*}_{\lambda}$ for all units. This simplification enables a transparent analysis of the estimator’s asymptotic variance and distribution.
Substituting $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ with its probability limit $\boldsymbol{\beta}^{*}_{\lambda}$ yields the infeasible oracle estimator
$\widehat{\tau}^{*}_{\textup{LDM}}=\frac{1}{n_{T}}\sum_{i:d_{i}=1}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big)-\frac{1}{n_{C}}\sum_{i:d_{i}=0}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big).$
Because treatment is assigned without replacement under complete random assignment, the oracle estimator can be viewed as a contrast of two finite-population averages drawn by sampling without replacement. Applying Hájek’s finite-population central limit theorem (hajek1960limiting) yields the following result.
@@
Under Assumptions˜1, 2 and 3, $\sqrt{n}\,(\widehat{\tau}_{\textup{LDM}}-\tau)$ converges in distribution to a Gaussian random variable with mean zero and variance
$\lim_{n\to\infty}\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}^{*}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2},$
@@
${\boldsymbol{\mu}}=\sqrt{\frac{n_{C}}{n_{T}}}\,{\bm{y}}^{(1)}+\sqrt{\frac{n_{T}}{n_{C}}}\,{\bm{y}}^{(0)}.$
The variance expression above parallels that of the LOORA-HT estimator but differs in two key aspects. First, because the treatment indicators are dependent under complete random assignment, the centering of both $\bm{X}$} and $\boldsymbol{\mu}$} appears explicitly in the limiting variance. Second, the scaling of $\boldsymbol{\mu}$} reflects the finite-sample proportions of treated and control units, corresponding to the population fractions $n_{T}/n$ and $n_{C}/n$.
Together, Lemma 6 and Theorem 7 show that LOORA-DM achieves an asymptotically normal distribution with a variance determined by the projection of the population signal onto the null space of the covariate space. This characterization implies the asymptotic efficiency of LOORA-DM similar to LOORA-HT, see Section˜3.4.
4.4 Confidence Intervals for LOORA-DM
We now turn to inference for the LOORA-DM estimator. As in the case of LOORA-HT, we construct confidence intervals by viewing the estimator as the second coefficient in an auxiliary regression and estimating its variance using the heteroskedasticity-robust (Huber–White) variance formula. This approach treats LOORA-DM as a two-step estimator: the first step computes regression-adjusted residuals, and the second step averages these residuals across treatment and control groups.
Regression representation. Recall that the LOORA-DM estimator can be expressed as
$\widehat{\tau}_{\textup{LDM}}$ $=\frac{1}{n_{T}}\sum_{i:d_{i}=1}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\big)-\frac{1}{n_{C}}\sum_{i:d_{i}=0}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\big).$
Define the adjusted residual vector and design matrix
${\bm{u}}:=\left(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\right)_{i=1}^{n},\qquad\boldsymbol{N}:=\begin{bmatrix}\boldsymbol{1}&~~{\bm{d}}\end{bmatrix}$,
where ${\bm{d}}=(d_{1},\ldots,d_{n})^{\top}$ is the treatment indicator vector. We now show that $\widehat{\tau}_{\textup{LDM}}$ equals the second coefficient from the least-squares projection of $\bm{u}$} onto $\boldsymbol{N}$.
Note that
$\boldsymbol{N}^{\top}\boldsymbol{N}=\begin{bmatrix}n&n_{T}\\ n_{T}&n_{T}\end{bmatrix},\qquad(\boldsymbol{N}^{\top}\boldsymbol{N})^{-1}=\frac{1}{n_{C}n_{T}}\begin{bmatrix}n_{T}&-n_{T}\\[1.99997pt] -n_{T}&n\end{bmatrix}$,
and
$\boldsymbol{N}^{\top}{\bm{u}}=\begin{bmatrix}\displaystyle\sum_{i=1}^{n}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\\[11.99998pt] \displaystyle\sum_{i:d_{i}=1}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\end{bmatrix}$.
Therefore,
$\big((\boldsymbol{N}^{\top}\boldsymbol{N})^{-1}\boldsymbol{N}^{\top}{\bm{u}}\big)_{2}$ $=-\frac{1}{n_{C}}\sum_{i=1}^{n}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}+\frac{n}{n_{C}n_{T}}\sum_{i:d_{i}=1}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$
$=\frac{1}{n_{T}}\sum_{i:d_{i}=1}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}-\frac{1}{n_{C}}\sum_{i:d_{i}=0}y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}.$
Thus $\widehat{\tau}_{\textup{LDM}}$ is precisely the second coefficient in the regression of $\bm{u}$} onto $\boldsymbol{N}$.
Huber–White variance estimator. Let
$\widehat{{\bm{V}}}_{\textup{HW}}=(\boldsymbol{N}^{\top}\boldsymbol{N})^{-1}\left(\boldsymbol{N}^{\top}\widehat{{\bm{S}}}\boldsymbol{N}\right)(\boldsymbol{N}^{\top}\boldsymbol{N})^{-1},\quad\text{where}\quad\widehat{{\bm{S}}}=\operatorname{diag}(\widehat{r}_{1}^{2},\ldots,\widehat{r}_{n}^{2}),$
and $\widehat{r}_{i}$ are the residuals from the regression,
@@ 4.4 Confidence Intervals for LOORA-DM
$\widehat{r}_{i}=\frac{y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}}{1-\widetilde{h}_{\lambda ii}}-\widehat{\theta}-\widehat{\tau}_{\textup{LDM}}\,d_{i},$
with $(\widehat{\theta},\widehat{\tau}_{\textup{LDM}})$ the estimated coefficients from regressing $\bm{u}$} on $\boldsymbol{N}$. Then the estimated variance of $\widehat{\tau}_{\textup{LDM}}$ is given by the $(2,2)$ entry of $\widehat{{\bm{V}}}_{\textup{HW}}$:
$\widehat{{\rm Var}}(\widehat{\tau}_{\textup{LDM}})=e_{2}^{\top}\widehat{{\bm{V}}}_{\textup{HW}}e_{2},\qquad e_{2}=(0,1)^{\top}.$
This estimator coincides with the standard heteroskedasticity-robust (Huber–White) variance estimator for the slope coefficient in an OLS regression.
Inference and consistency. The regression formulation above implies that valid confidence intervals for the average treatment effect can be constructed using the estimated standard error
$\widehat{\sigma}=\sqrt{e_{2}^{\top}\widehat{{\bm{V}}}_{\textup{HW}}e_{2}},$
and normal critical values. Analogous to the case of LOORA-HT, the consistency of $\widehat{\tau}_{\textup{LDM}}$ does not depend on the consistency of $\widehat{\boldsymbol{\beta}}_{\lambda}$ itself: even if the first-stage ridge estimates are biased, the two-step estimator remains consistent for the average treatment effect under complete random assignment.
@@ 5 Discussion
The estimators developed in this paper contribute to a long line of work on regression adjustment and design-based inference in randomized experiments. A central goal of this literature is to reconcile two desiderata that are often in tension: finite-sample unbiasedness under the experimental design, and efficiency gains through regression-based covariate adjustment. Classical estimators such as the Horvitz–Thompson and difference-in-means achieve the former but not the latter, while regression-adjusted estimators improve efficiency at the cost of bias or instability in small samples. The LOORA estimators bridge this gap by providing finite-sample unbiased and asymptotically efficient estimators that remain robust to high-leverage covariates.
To highlight how the LOORA estimators relate to existing methods, Table 1 summarizes the key properties of several classical estimators under simple and complete randomization. The comparison focuses on unbiasedness, asymptotic efficiency relative to the benchmark of lin2013agnostic, and robustness to leverage. The table emphasizes that while classical regression adjustments can reduce variance, none of them simultaneously ensure all three properties.
@@
Classical estimators. Under simple randomization, the Horvitz–Thompson (HT) estimator is exactly unbiased but can exhibit high variance. Regression adjustment aims to improve efficiency by controlling for observed covariates. The most commonly used adjusted estimator, $\widehat{\tau}_{\text{ADJ}}$, regresses $\bm{y}$} on $\bm{X}$} and the treatment indicator $\bm{d}$} with intercept, producing a consistent estimator only when treatment effects are homogeneous. The interacted version, $\widehat{\tau}_{\text{INT}}$, adds treatment-by-covariate interactions and is asymptotically efficient under both simple and complete randomization (lin2013agnostic). However, both estimators can become unstable when the covariate matrix contains influential or high-leverage observations, and neither is finitely unbiased under the randomization design (young2019channeling).
@@
Leave-one-out regression adjustment. The LOORA framework modifies regression adjustment to preserve finite-sample unbiasedness without sacrificing efficiency. By using leave-one-out fitted values, the LOORA-HT and LOORA-DM estimators remove the bias introduced by in-sample residual reuse while maintaining asymptotic optimality. Under simple randomization, LOORA-HT achieves the same variance bound as the lin2013agnostic estimator but is exactly unbiased and robust to leverage. Under complete randomization, LOORA-DM extends these results to dependent assignment structures by appropriately rescaling the leave-one-out predictions. Both estimators therefore combine the finite-sample exactness of HT and DM with the asymptotic efficiency of regression adjustment.
@@
Relation to recent work. Recent contributions have revisited the foundations of design-based inference from distinct perspectives. ghadiri2023finite develop a finite-sample unbiased regression-adjusted estimator under simple random assignment, using a cross-fitted formulation that motivates the present work. However, their approach does not attain the asymptotically efficient variance bound. harshaw2022design, in contrast, take a design-based perspective: rather than modifying the estimator, they optimize the randomization mechanism itself to minimize the variance of the unadjusted Horvitz–Thompson estimator. Finally, spiess2025optimal characterize the necessary conditions under which design-based estimators such as Horvitz–Thompson or difference-in-means are unbiased. For the HT estimator, these conditions require a leave-one-out structure, which our LOORA-HT estimator satisfies by construction. For the DM estimator, unbiasedness requires a leave-two-out structure; we show that LOORA-DM can be equivalently viewed as a leave-two-out estimator, thereby satisfying Spiess’s condition. Hence, our estimators are not only consistent with the theoretical framework of spiess2025optimal, but also provide constructive, closed-form realizations of unbiased estimators that satisfy these necessary design-based criteria. Detailed comparisons with these works are presented in the subsections that follow.
Summary of properties. As shown in Table 1, the LOORA-HT and LOORA-DM estimators uniquely satisfy all three desirable properties: they are finitely unbiased, asymptotically efficient, and robust to high-leverage covariate realizations under their respective randomization schemes. Classical regression-adjusted estimators achieve at most two of these properties. By contrast, the LOORA estimators combine the robustness of design-based estimators with the efficiency of regression adjustment in a single, tractable framework.
@@
unbiased attains lin2013agnostic robust to
asymptotic high
variance leverage
@@
$\widehat{\tau}_{\text{ADJ}}$ (Classic Regression Adjustment)
$\widehat{\tau}_{\text{INT}}$ (Interacted Regression Adjustment lin2013agnostic)
$\widehat{\tau}_{\text{LOORA-DM}}$
@@ 5.1 Comparison with Cross-Fitted Regression-Adjusted HT
ghadiri2023finite propose a finite-sample unbiased regression-adjusted Horvitz–Thompson estimator based on a cross-fitting approach; that is, the sample is split into two groups, a regression vector is learned separately within each group, and the fitted vector from one group is used to adjust the outcomes of the other. They consider only the setting where $p_{i}=0.5$ for all $i$, and provide the following upper bound on the variance of their estimator:
$\frac{8(1+\varepsilon)}{n^{2}}\min_{{\bm{b}}\in\mathbb{R}^{d}}\Big(\left\|{\bm{X}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big)+\frac{32d}{n^{2}}\left\|{\bm{y}}^{(1)}-{\bm{y}}^{(0)}\right\|_{\infty}^{2}.$ (17)
By Lemma˜2, with their choice of regularization parameter $\lambda=100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}$, the leverage scores satisfy $h_{i}({\bm{X}},\lambda)\leq\frac{1}{1+100\log(n/\delta)}$. Moreover, since all $p_{i}=0.5$, it follows that
$\max_{i\in[n]}\left\|\widetilde{{\bm{x}}}_{i}\right\|_{2}=2\max_{i\in[n]}\left\|{\bm{x}}_{i}\right\|_{2}.$
Hence, noting that $4\lambda=100\log(n/\delta)\,\zeta_{\widetilde{{\bm{X}}}}^{2}$, we have $h_{i}({\bm{X}},\lambda)=h_{i}(\widetilde{{\bm{X}}},4\lambda)$.
Now consider
$\boldsymbol{\beta}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{d}}\Big(\left\|{\bm{X}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big)$
$=\frac{1}{2}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{d}}\Big(\left\|\widetilde{{\bm{X}}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{\widetilde{{\bm{X}}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big).$
The variance of our estimator, from Theorem˜1, is bounded by
$\frac{1}{n^{2}(1-\frac{1}{101})^{2}}\sum_{i=1}^{n}(\widetilde{{\bm{x}}}_{i}^{\top}(2\boldsymbol{\beta})-\mu_{i})^{2}+\frac{8d}{n^{2}}\left\|{\bm{t}}/{\bm{r}}\right\|_{\infty}^{2}$
$=\frac{1.0201}{n^{2}}\sum_{i=1}^{n}({\bm{x}}_{i}^{\top}\boldsymbol{\beta}-\mu_{i})^{2}+\frac{8d}{n^{2}}\left\|{\bm{y}}^{(1)}-{\bm{y}}^{(0)}\right\|_{\infty}^{2}.$
Relative to the bound of ghadiri2023finite in (17), the bound above improves the first term by approximately a factor of eight. It also eliminates the additional ridge regularization penalty term appearing in (17), and improves the final term by a factor of four.
Finally, it is worth noting that ghadiri2023finite do not provide an inference procedure or a method for constructing confidence intervals, whereas we provide a two-step estimator perspective that allows consistent variance estimation and valid confidence interval construction.

### v2

@@ 4.3 Asymptotic Normality of LOORA-DM
This section analyzes the large-sample behavior of LOORA-DM under complete random assignment. It shows that, in large samples, LOORA-DM satisfies a central limit theorem. Similar to Section˜3.3, we study the behavior of truncated sequences $({\bm{x}}_{i},y_{i}^{(1)},y_{i}^{(0)})_{i\in[n]}$ as $n\to\infty$, and assume that the ratio $n_{T}/n$ is bounded away from zero and one; see Assumption˜2.
@@
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{k}}\big\|{\bm{X}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}^{(-i)}_{-i}\big\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2}.$
To study the limiting behavior of these regression adjustments, define the following population version of $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$
$\boldsymbol{\beta}_{\lambda}=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in{\mathbb{R}}^{k}}\left\|{\bm{X}}{\bm{b}}-\widetilde{{\boldsymbol{\mu}}}\right\|_{2}^{2}+\lambda\|{\bm{b}}\|_{2}^{2},$ (11)
$\boldsymbol{\beta}_{\lambda}$ represents the deterministic coefficient vector that best captures the relation between covariates and the signal component $\widetilde{{\boldsymbol{\mu}}}$ in the population.
@@
Suppose Assumptions˜1, 2 and 3 hold. Let $\lambda\geq 0$ be fixed. If
$\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}\allowbreak-\allowbreak(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}\to\infty,$
then
$\frac{\sqrt{n}({\widehat{\tau}}_{\textup{LDM}}-\tau)}{\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}/\sqrt{n}}$
converges in distribution to a standard normal random variable.
The variance expression above parallels that of the LOORA-HT estimator but differs in two key aspects. First, the large sample variance of the difference-in-means estimator under complete random assignment automatically recenters $\bm{X}$} and $\widetilde{{\boldsymbol{\mu}}}$. In contrast, LOORA-HT requires the inclusion of an intercept term to obtain a similar variance expression. Second, the scaling of $\widetilde{{\boldsymbol{\mu}}}$ reflects the finite-sample proportions of treated and control units (equation (8)).
Theorem 6 shows that LOORA-DM achieves an asymptotically normal distribution with a variance determined by the projection of the population signal onto the null space of the covariate matrix. It implies that LOORA-DM is as efficient in large samples as the interacted adjustment estimator of Lin (2013).
4.4 Confidence Intervals for LOORA-DM
The next theorem provides a consistency result for an estimator of the variance of LOORA-DM.
Theorem 7.
Suppose Assumptions˜1, 2 and 3 hold. Let $\lambda\geq 0$ be fixed, define
$\widehat{V}_{\textup{LDM}}=\frac{n-1}{n_{C}(n_{C}-1)}\sum_{i\in[n]:d_{i}=0}(\widehat{s}_{i}^{(0)})^{2}+\frac{n-1}{n_{T}(n_{T}-1)}\sum_{i\in[n]:d_{i}=1}(\widehat{s}_{i}^{(1)})^{2},$
@@ 4.4 Confidence Intervals for LOORA-DM
$\widehat{s}_{i}^{(0)}$ $=y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}-\frac{1}{n_{C}}\sum_{j\in[n]:d_{j}=0}y_{j}-{\bm{x}}_{j}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-j)},$
$\widehat{s}_{i}^{(1)}$ $=y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}-\frac{1}{n_{T}}\sum_{j\in[n]:d_{j}=1}y_{j}-{\bm{x}}_{j}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-j)}.$
Moreover let
@@
$W_{n}=\frac{1}{n}\left(\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{\lambda}-(\widetilde{{\boldsymbol{\mu}}}-\overline{\widetilde{{\boldsymbol{\mu}}}})\right\|_{2}^{2}+\sum_{i=1}^{n}(\tau_{i}-\tau)^{2}\right).$
If $W_{n}$ is bounded away from zero, then
$\frac{\widehat{V}_{\textup{LDM}}}{W_{n}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}1.$
As with LOORA-HT, and more generally with finite-population variance estimation for treatment-effect estimators, $\widehat{V}_{\textup{LDM}}$ is consistent under treatment-effect homogeneity and conservative otherwise. Large-sample confidence intervals for $\tau$ follow by combining this variance estimator with quantiles of the standard normal distribution.
@@ 5 Discussion
LOORA targets three properties that existing estimators do not achieve simultaneously: finite-sample unbiasedness, asymptotic efficiency, and robustness to high-leverage observations. Classical estimators such as Horvitz–Thompson and the difference-in-means are exactly unbiased, but they leave efficiency gains on the table. Regression-adjusted estimators recover those gains asymptotically, yet they introduce finite-sample bias and can become unstable in the presence of high-leverage observations. LOORA delivers all three properties simultaneously: it preserves exact unbiasedness, attains the Lin (2013) efficiency bound in large samples, and remains stable under influential covariates through its leave-one-out construction.
Table 1 summarizes the properties for LOORA and competing estimators under simple and complete randomization.
@@
Classical estimators. Under simple randomization, the Horvitz–Thompson (HT) estimator is exactly unbiased but can exhibit high variance. Regression adjustment aims to improve efficiency by controlling for observed covariates. The most commonly used adjusted estimator, $\widehat{\tau}_{\text{ADJ}}$, regresses $\bm{y}$} on $\bm{X}$} and the treatment indicator $\bm{d}$} with intercept, producing a consistent estimator only when treatment effects are homogeneous. The interactive version, $\widehat{\tau}_{\text{INT}}$, adds treatment-by-covariate interactions and is asymptotically efficient under both simple (with equal treatment probability assignment for all units) and complete randomization (Lin, 2013). However, both estimators, $\widehat{\tau}_{\text{ADJ}}$ and $\widehat{\tau}_{\text{INT}}$, can become unstable when the sample contains high-leverage observations, and neither is unbiased.
@@
Leave-one-out regression adjustment. The LOORA framework modifies regression adjustment to preserve finite-sample unbiasedness without sacrificing efficiency. By using leave-one-out fitted values, the LOORA-HT and LOORA-DM estimators remove the bias induced by regression adjustment while maintaining asymptotic optimality. Under simple randomization, LOORA-HT achieves the same variance bound as the Lin (2013) estimator but is exactly unbiased and robust to leverage. Under complete randomization, LOORA-DM extends these results to dependent assignment structures by appropriately rescaling the leave-one-out predictions. Both estimators combine the finite-sample exactness of HT and DM with the asymptotic efficiency of regression adjustment.
@@
Relation to recent work. Recent work has proposed innovative methods for covariate adjustment in randomized experiments. Ghadiri et al. (2023) develops a cross-fitting unbiased regression-adjusted estimator for simple random assignment. This approach, however, does not attain the Lin (2013) variance in large samples. Harshaw et al. (2024), in contrast, takes a design-based perspective: rather than modifying the estimator, they optimize the randomization mechanism itself to minimize the variance of the unadjusted Horvitz–Thompson estimator. Finally, Spiess (2025) characterizes the necessary conditions under which design-based estimators such as Horvitz–Thompson or difference-in-means are unbiased. For the HT estimator, these conditions require a leave-one-out structure, which our LOORA-HT estimator satisfies by construction. For the DM estimator, unbiasedness requires a leave-two-out structure. We show below that LOORA-DM can be represented as a leave-two-out estimator, thereby satisfying Spiess’s condition. Hence, our estimators are not only consistent with the theoretical framework of Spiess (2025), but also provide constructive, closed-form realizations of unbiased estimators that satisfy these necessary design-based criteria. The remainder of this section provides a detailed comparison with prior work.
@@
unbiased attains Lin (2013) robust to
asymptotic variance high leverage
@@
     $\widehat{\tau}_{\text{ADJ}}$ (Linear regression adjustment)
     $\widehat{\tau}_{\text{INT}}$  (Interactive regression adjustment, Lin (2013))
     $\widehat{\tau}_{\text{LDM}}$
@@ 5.1 Comparison with Cross-Fitted Regression-Adjusted HT
Ghadiri et al. (2023) proposes a finite-sample unbiased regression-adjusted Horvitz–Thomp-son estimator built on cross-fitting. The procedure splits the sample into two groups, learns a regression vector within each group, and then uses each group’s fitted vector to adjust the outcomes in the other. Ghadiri et al. (2023) considers only the case of constant $p_{i}=0.5$ for all $i$ and derives a high-probability upper bound on the variance of their estimator. Specifically, conditional on an event $\mathcal{E}$ that holds with probability at least $1-\delta$, the variance of the estimator in Ghadiri et al. (2023) is bounded by
$\frac{8(1+\varepsilon)}{n^{2}}\min_{{\bm{b}}\in\mathbb{R}^{k}}\Big(\left\|{\bm{X}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big)+\frac{32k}{n^{2}}\left\|{\bm{y}}^{(1)}-{\bm{y}}^{(0)}\right\|_{\infty}^{2},$ (12)
[…]
$\boldsymbol{\beta}$ $=\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\Big(\left\|{\bm{X}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big).$
Note that for any ${\bm{b}}\in{\mathbb{R}}^{k}$,
$\left\|{\bm{X}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{{\bm{X}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}=\left\|\widetilde{{\bm{X}}}({\bm{b}}/2)-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{\widetilde{{\bm{X}}}}^{2}\left\|{\bm{b}}/2\right\|_{2}^{2}.$
Therefore,
$\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}\in\mathbb{R}^{k}}\Big(\left\|\widetilde{{\bm{X}}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+100\log(n/\delta)\,\zeta_{\widetilde{{\bm{X}}}}^{2}\left\|{\bm{b}}\right\|_{2}^{2}\Big)=\boldsymbol{\beta}/2.$
For $\lambda=100\log(n/\delta)\,\zeta_{\widetilde{{\bm{X}}}}^{2}$ with $\delta\leq\exp(-1)$, the variance of LOORA-HT from Theorem˜1, is bounded by
$\left(1-\frac{1}{101}\right)^{-2}\left(\frac{1}{n^{2}}\sum_{i=1}^{n}(\widetilde{{\bm{x}}}_{i}^{\top}(\boldsymbol{\beta}/2)-\mu_{i})^{2}+\frac{2k}{n^{2}}\left\|{\bm{t}}/{\bm{r}}\right\|_{\infty}^{2}\right)$
$=\left(\frac{101}{100}\right)^{2}\left(\frac{1}{n^{2}}\sum_{i=1}^{n}({\bm{x}}_{i}^{\top}\boldsymbol{\beta}-\mu_{i})^{2}+\frac{2k}{n^{2}}\left\|{\bm{y}}^{(1)}-{\bm{y}}^{(0)}\right\|_{\infty}^{2}\right).$
Relative to the bound of Ghadiri et al. (2023) in (12), the bound above improves the first term by roughly a factor of eight. It also eliminates the additional ridge-regularization penalty term that appears in (12) and improves the final term by about a factor of sixteen.
In addition, unlike the procedure in Ghadiri et al. (2023), LOORA comes with consistent variance estimation and valid confidence intervals.